Find the angle θ. (Use either the Law of Sines or the Law of Cosines, as appropriate.) (Round your answer to one decimal place.)
a = 14, c = 8, C = 30°

Answer by stanbon(75887)

Find the angle θ. (Use either the Law of Sines or the Law of Cosines, as appropriate.) (Round your answer to one decimal place.)
a = 14, c = 8, C = 30°
---
sin(A)/a = sin(C)/c
---
sin(A) = (a/c)*sin(C)
---
sin(A) = (14/8)*sin(30)
---
sin(A) = 14/8*1/2 = 7/8
---
A = sin^-1(7/8) = 61.045 degrees
===================================
Cheers,
Stan H.
=================